If the demand function faced by a firm is:

Q = 90 – 2P

TC = 2 + 57Q – 8Q2 + Q3


Determine the level of output at which the firm maximizes the profit.

Determine the best level of output for the above question by the MR and MC approach.



1
Expert's answer
2021-05-06T13:15:09-0400

"Q = 90 - 2P \\

P = 45 - 0.5Q \\

TC = 2 + 57Q - 8Q^2 + Q^3"

The firm’s profit-maximizing output level is:

"MR = MC \\

MR = TR'(Q) = 45 -Q \\

MC = TC'(Q) = 3Q^2 - 16Q + 57 \\

3Q^2 - 16Q + 57 = 45 - Q \\

3Q^2 - 15Q + 12 = 0 \\

Q^2 - 5Q + 4 = 0 \\

Q_1 = 4 \; units \\

Q_2 = 1\; unit"